Graph the expression sec2 x - sec2 x sin2 x on your calculator. Determine what constant or single function is equivalent to the given expression.
Answer by lwsshak3(11628)

Graph the expression sec2 x - sec2 x sin2 x on your calculator. Determine what constant or single function is equivalent to the given expression.
**
sec2 x - sec2 x sin2 x
=1/cos^2x-1/cos^2x*sin^2x
=1/cos^2x-sin^2x/cos^2x
=(1-sin^2x)/cos^2x
=cos^2x/cos^2x
=1